# Index Terms:

### v1

@@ Abstract
Downlink MIMO Channel Estimation from Bits: Recoverability and Algorithm Rajesh Shrestha, Mingjie Shao, Mingyi Hong, Wing-Kin Ma, and Xiao Fu R. Shrestha and M. Shao contributed equally. R. Shrestha and X. Fu are with the School of EECS at Oregon State University, Corvallis, OR, USA. M. Shao is with the School of Information Science and Engineering, Shandong University, Qingdao, China. W.-K. Ma is with the Department of EE, The Chinese University of Hong Kong, Hong Kong SAR of China. M. Hong is with the Department of ECE at University of Minnesota, Minneapolis, MN, USA. The work of R. Shrestha and X. Fu was supported in part by the National Science Foundation (NSF) under project NSF CCF-2210004.
@@ I Introduction
Vector quantization (VQ) codebook-based schemes [3, 2] are arguably the most well-known methods for limited feedback, but naive VQ encounters scalability challenges in the era of massive MIMO. More recent works such as [9] and [10] considered more advanced precoding matrix indicator (PMI) and channel quality indicator (CQI) feedback, but similar challenges remain. Many works use low-dimensional parameterizations of the large channel matrix (e.g., using stochastic models [11, 12] or geometric models [13, 5, 14]), and estimate the key parameters of the channel at the UE [15, 16, 17, 18, 19]. This way, the UE can just feedback the key parameters other than the entire channel matrix, substantially reducing the overhead. In particular, the works [20, 18, 5, 16, 21, 14] employ the double directional parameterization, where the channel matrix is represented by $K$ paths with path losses, angle-of-departure (AoD) and angle-of-arrivals (AoA) being the parameters. The path losses, AoDs and AoDs are then estimated via methods such as tensor decomposition [15, 19, 22] and compressive sensing [18, 23, 15, 14]. A notable challenge in this line of work is that the UEs have limited computational resources (e.g., power, memory, and CPU capacity), yet channel parameter estimation is often a nontrivial task. As a workaround, the works [15, 5, 14] advocated schemes where the UE only compresses the channel information, while the major computations for CSI recovery are conducted at the BS. Two dictionary-based sparse algorithms were proposed in [15, 14], using discretized spatial representations of AoD and AoD to construct the dictionary.
@@ II-A Signal Model
where $\beta_{k}\in\mathbb{C}$ is the channel path loss of the $k$th path; $\theta_{k},\phi_{k}\in[-\pi,\pi]$ are the angle of arrival (AoA) and angle of departure (AoD) of the $k$th path, respectively; ${\bm{a}}_{r}(\theta_{k})$ and ${\bm{a}}_{t}(\phi_{k})$ are the steering vectors associated with the AoA $\theta_{k}$ and AoD $\phi_{k}$, respectively, which are given by
@@ III-A1 Operations at UE
In this scheme, we assume that the UE first estimates $\bm{H}$ via relatively simple operations, e.g., least squares. In practice, one can choose the pilot signal ${\bm{S}}$ to be row orthogonal, i.e., ${\bm{S}}{\bm{S}}^{\rm{H}}={\bm{I}}.$ As a result, the LS estimation of ${\bm{H}}$ under (1) can be efficiently computed via
@@
$\widehat{{\bm{h}}}=\mbox{vec}(\overline{{\bm{H}}}),\quad\overline{{\bm{H}}}=\begin{bmatrix}\Re(\widehat{{\bm{H}}})\\ \Im(\widehat{{\bm{H}}})\end{bmatrix},$ (6)
where $\Re(\widehat{{\bm{H}}})$ and $\Im(\widehat{{\bm{H}}})$ represent the real and imaginary parts of $\hat{{\bm{H}}}$, respectively. We denote $\widehat{{\bm{h}}}\in\mathbb{R}^{J}$ as the vectorized $\widehat{{\bm{H}}}$ with $J=2MN$. We compress the channel $\hat{{\bm{h}}}$ by a random matrix ${\bm{A}}\in\mathbb{R}^{R\times J}$ with $R\ll J$ and
@@
The compression matrix ${\bm{A}}$ is known to both the BS and the UE. The compressed channel is given by
@@ III-A2 MLE Recovery at BS
The BS receives the compressed and quantized CSI ${\bm{r}}$ through a designated signaling channel. Subsequently, the BS estimates the full channel matrix ${\bm{H}}$ from ${\bm{r}}$. Based on the compression and quantization process in (5)-(8), the likelihood function $p({\bm{r}}|{\bm{h}})$ is given by
$\begin{split}p({\bm{r}}|{\bm{h}})=&\leavevmode\nobreak\ \prod_{i=1}^{R}p(r_{i}|{\bm{h}}),\\ p(r_{i}|{\bm{h}})=&\leavevmode\nobreak\ \Phi\left(\frac{\bar{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}}{\sigma_{v}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}}{\sigma_{v}}\right),\end{split}$ (9)
@@
As a result, the MLE of ${\bm{h}}$ can be expressed as follows:
$\begin{split}\operatorname*{\textrm{minimize}}_{{\bm{z}},{\bm{H}}}&\leavevmode\nobreak\ -\log p({\bm{r}}|{\bm{h}})\\ \mbox{subject\leavevmode\nobreak\ to}&\leavevmode\nobreak\ {\cal G}({\bm{z}})={\bm{H}},\\ \end{split}$ (10)
where we stack the channel parameters ${\bm{z}}=(\bm{\theta},\bm{\phi},\bm{\beta})$, and the operator $\cal G$ denotes the channel model (4) that maps the channel parameters ${\bm{z}}$ to the channel ${\bm{H}}$
@@ III-B1 Operations at UE
where $v_{t}\sim{\cal N}(0,\sigma^{2}_{v})$. Then, the UE sends $r_{t}$ for $t=1,\ldots,T$. Compared to Scheme 1, we no longer need to estimate the channel by the LS at the UE side.
@@ III-B2 MLE Recovery at BS
Note that $\tilde{{\bm{y}}}_{t}$ in (11) can be written as
@@
$\begin{split}{\bm{D}}=\begin{bmatrix}\tilde{{\bm{s}}}_{1}^{\top}\otimes{\bm{a}}_{1}^{\top}\\ \vdots\\ \tilde{{\bm{s}}}_{T}^{\top}\otimes{\bm{a}}_{T}^{\top}\end{bmatrix},\quad\check{{\bm{h}}}=\mbox{vec}(\tilde{{\bm{H}}}),\leavevmode\nobreak\ \check{{\bm{n}}}=\begin{bmatrix}{\bm{a}}_{1}^{\top}\tilde{{\bm{n}}}_{1}\\ \vdots\\ {\bm{a}}_{T}^{\top}\tilde{{\bm{n}}}_{T}\end{bmatrix},\end{split}$
and we have used ${\bm{a}}_{i}^{\top}\tilde{{\bm{H}}}\tilde{{\bm{s}}}_{i}=(\tilde{{\bm{s}}}_{i}^{\top}\otimes{\bm{a}}_{i}^{\top})\mbox{vec}(\tilde{{\bm{H}}})$ [34]. Then, the feedback signal model can be rewritten as
@@
where $\check{{\bm{v}}}=\check{{\bm{n}}}+{\bm{v}}$ and it can be shown that $\check{{\bm{v}}}\sim{\cal N}(\bm{0},\check{\sigma}^{2}{\bm{I}})$ for $\check{\sigma}^{2}=\sigma_{n}^{2}+\sigma_{v}^{2}$. Similar to Scheme 1, the BS estimates $\check{{\bm{h}}}$ through the following ML estimation problem
$\begin{split}\operatorname*{\textrm{minimize}}_{{\bm{z}},{\bm{H}}}&\leavevmode\nobreak\ -\log p({\bm{r}}|\check{{\bm{h}}})\\ \mbox{subject\leavevmode\nobreak\ to}&\leavevmode\nobreak\ {\cal G}({\bm{z}})={\bm{H}},\\ \end{split}$ (15)
@@
$\begin{split}p({\bm{r}}|\check{{\bm{h}}})=&\leavevmode\nobreak\ \prod_{i=1}^{N_{S}}p(r_{i}|\check{{\bm{h}}})\\ p(r_{i}|\check{{\bm{h}}})=&\leavevmode\nobreak\ \Phi\left(\frac{\bar{b}[r_{i}]-{\bm{d}}_{i}^{\top}\check{{\bm{h}}}}{\check{\sigma_{v}}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{d}}_{i}^{\top}\check{{\bm{h}}}}{\check{\sigma_{v}}}\right),\end{split}$
@@
In general, Scheme 1 has more operations performed by the UE as it involves estimating ${\bm{H}}$, compression, and quantization. These give rise to $O(MNT+MNR)$ flops. Scheme 2 saves time at UE as it directly compresses the ${\bm{H}}\bm{S}$ without extracting ${\bm{H}}$ first, which only uses $O(MT)$ flops. Note that the flop reduction at the UE by Scheme 2 can be substantial relative to Scheme 1, as $N$ could be (much) larger than $M$. Nonetheless, as one will see in the next section, Scheme 2 gains in efficiency but loses in sample complexity—that is, Scheme 2 requires slightly more feedback bits to attain the same CSI recovery accuracy of Scheme 1.
@@ IV-A Preliminaries
In the channel model (2), it is reasonable to assume that the channel path loss $\beta_{k}$ is bounded. We denote the maximum amplitude of $\beta_{k}$ as $\kappa$, i.e., $|\beta_{k}|\leq\kappa$ for all $i=1,\ldots,K$. The parameter space of ${\bm{z}}$, denoted by $\mathcal{Z}$, is then described by
@@
that is, the function ${\cal G}$ is Lipstchiz continuous.
Note that $L_{\cal G}<\infty$ exists because the array manifolds and the double directional channel model do not give rise to erratic changes under continuous inputs $\bm{z}$; see Appendix D in the supplementary material of this paper.
@@
Note that when $x$ is bounded, the negative log-likelihood function satisfies
@@
for a certain constant $U_{f}>0$. In addition, as the gradient of $f_{r}(x)$ is continuous, by the extreme value theorem we have
@@
for some constant $F_{f}>0$. The constant $F_{f}$ being bounded away from zero means that the function $f_{r}(x)$ does not change too slowly.
@@ IV-B Recoverability Analysis of Scheme 1
Assume that the entries of ${\bm{A}}\in\mathbb{R}^{R\times J}$ satisfy $a_{i,j}\sim{\cal N}(0,1/R)$ for all $i,j$. In addition, assume that
@@
$\begin{split}&\leavevmode\nobreak\ \frac{\|{\bm{h}}^{\natural}-{\bm{h}}^{\star}\|_{2}^{2}}{MN}\\ \leq&\leavevmode\nobreak\ \frac{64}{\sqrt{R}}\left(4\frac{L_{f}}{F_{f}}(\tau+\frac{3}{4}\nu)+9\frac{U_{f}}{F_{f}}\sqrt{\log(2/\eta)}\right)+\frac{32\psi^{2}}{R}\\ \end{split}$
with probability at least $1-3\eta-e^{-\Omega(R)}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{R}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}$. and $\psi=\nicefrac{{1}}{{K}}+\nicefrac{{\nu}}{{2}}+2\sqrt{2}+\sqrt{\nicefrac{{2J}}{{R}}}$.
@@ IV-C Recoverability Analysis of Scheme 2
Assume that ${\bm{D}}$ has a size of $T\times J$ with $J=4MN$, where ${\bm{D}}=\nicefrac{{1}}{{T}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$ and the entries of $\bm{S}$ and $\bm{A}$ are independently drawn from the standard normal distribution. Assume that the following holds:
@@
$\begin{split}&\leavevmode\nobreak\ \frac{\|{\bm{h}}^{\natural}-{\bm{h}}^{\star}\|_{2}^{2}}{MN}\\ \leq&\leavevmode\nobreak\ \frac{64}{\sqrt{T}}\left(4\frac{L_{f}}{F_{f}}(\tau+\frac{3}{4}\nu)+9\frac{U_{f}}{F_{f}}\sqrt{\log(2/\eta)}\right)+\frac{32\psi^{2}}{T}\end{split}$
with probability at least $1-3\eta-e^{-\mathbf{\Omega}(\sqrt{T})}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{T}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}$ and $\psi=\nicefrac{{1}}{{K}}+\nicefrac{{\nu}}{{2}}+2\sqrt{2}+\sqrt{\nicefrac{{2J}}{{T}}}$.
The proof of Theorem 2 is relegated to Appendix E in the supplementary material of this paper.

### v2

@@ Abstract
Downlink MIMO Channel Estimation from Bits: Recoverability and Algorithm Rajesh Shrestha, Mingjie Shao, Mingyi Hong, Wing-Kin Ma, and Xiao Fu R. Shrestha and M. Shao contributed equally. R. Shrestha and X. Fu are with the School of EECS at Oregon State University, Corvallis, OR 97331, USA. M. Shao was with the School of Information Science and Engineering, Shandong University, Qingdao 266237, china. He is now with the Key Laboratory of Systems and Control, Institute of Systems Science, Academy of Mathematics and Systems Science (AMSS), chinese Academy of Sciences (CAS), Beijing 1100149, China. W.-K. Ma is with the Department of EE, The Chinese University of Hong Kong, Hong Kong SAR of China. M. Hong is with the Department of ECE at University of Minnesota, Minneapolis, MN 55455, USA. (Corresponding author: Xiao Fu) The work of R. Shrestha and X. Fu was supported in part by the National Science Foundation (NSF) under project NSF CCF-2210004. The work of M. Hong was supported in part by the NSF CIF-2414372 and NSF ECCS-2426064. The work of W.-K. Ma was supported by a General Research Fund of Hong Kong Research Grant Council under Project ID 14203721. The work by M. Shao was supported in part by the National Natural Science Foundation of China under Grant 62401340 and the Natural Science Foundation of Shandong Province under Grant ZR2023QF103.
@@ I Introduction
Vector quantization (VQ) codebook-based schemes [3, 2] are arguably the most well-known methods for limited feedback, but naive VQ encounters scalability challenges in the era of massive MIMO. More recent works such as [9] and [10] considered more advanced precoding matrix indicator (PMI) and channel quality indicator (CQI) feedback, but similar challenges remain. Many works use low-dimensional parameterizations of the large channel matrix (e.g., using stochastic models [11, 12] or geometric models [13, 5, 14]), and estimate the key parameters of the channel at the UE [15, 16, 17, 18, 19]. This way, the UE can just feedback the key parameters other than the entire channel matrix, substantially reducing the overhead. In particular, the works [20, 18, 5, 16, 21, 14] employ the double directional parameterization, where the channel matrix is represented by $K$ paths with path losses, angle-of-departure (AoD) and angle-of-arrivals (AoA) being the parameters. The path losses, AoDs and AoAs are then estimated via methods such as tensor decomposition [15, 19, 22] and compressive sensing [18, 23, 15, 14]. A notable challenge in this line of work is that the UEs have limited computational resources (e.g., power, memory, and CPU capacity), yet channel parameter estimation is often a nontrivial task. As a workaround, the works [15, 5, 14] advocated schemes where the UE only compresses the channel information, while the major computations for CSI recovery are conducted at the BS. Two dictionary-based sparse algorithms were proposed in [15, 14], using discretized spatial representations of AoD and AoA to construct the dictionary.
@@ II-A Signal Model
where $\beta_{k}\in\mathbb{C}$ is the channel path loss of the $k$th path; $\theta_{k},\phi_{k}\in[-\pi,\pi]$ are AoA and AoD of the $k$th path, respectively; ${\bm{a}}_{r}(\theta_{k})$ and ${\bm{a}}_{t}(\phi_{k})$ are the steering vectors associated with the AoA $\theta_{k}$ and AoD $\phi_{k}$, respectively, which are given by
@@ III-A1 Operations at UE
In this scheme, we assume that the UE first estimates $\bm{H}$ via relatively simple operations, e.g., least squares.
In practice, one can choose the pilot signal $\bm{S}$} to be row orthogonal, i.e., ${\bm{S}}{\bm{S}}^{\rm{H}}={\bm{I}}.$ As a result, the LS estimation of $\bm{H}$} under (1) can be efficiently computed via
@@
[…]
@@
The compression matrix $\bm{A}$} is known to both the BS and the UE. In practice, $\bm{A}$} can be agreed upon the BS and UE via protocols; it can be generated using the same pseudo random number generator and seed at both sides before each transmission.
The compressed channel is given by
@@ III-A2 MLE Recovery at BS
The BS receives the compressed and quantized CSI $\bm{r}$} through a designated signaling channel. Subsequently, the BS estimates the full channel matrix $\bm{H}$} from $\bm{r}$}. Based on the compression and quantization process in (5)-(8), the likelihood function $p({\bm{r}}|{\bm{h}};{\bm{A}})$ is given by
$\begin{split}p({\bm{r}}|{\bm{h}};{\bm{A}})=&\leavevmode\nobreak\ \prod_{i=1}^{R}p(r_{i}|{\bm{h}};{\bm{a}}_{i}),\\ p(r_{i}|{\bm{h}};{\bm{a}}_{i})=&\leavevmode\nobreak\ \Phi\left(\frac{\bar{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}}{\sigma_{v}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}}{\sigma_{v}}\right),\end{split}$ (9)
where $\Phi(x)=\int_{-\infty}^{x}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2}\ dt$ is the cumulative distribution function of standard Gaussian distribution. This formulation arises from the dithering process, which induces a distribution $x_{i}\sim{\cal N}({\bm{a}}_{i}^{T}{\bm{h}},\sigma_{v}^{2})$. Consequently, the probability of the quantized value $r_{i}$ corresponds to the probability mass of this Gaussian distribution within the boundaries $(\underline{b}[r_{i}],\bar{b}[r_{i}])$, as expressed in (9).
As a result, the MLE of $\bm{h}$} can be expressed as follows:
@@
where we stack the channel parameters ${\bm{z}}=(\bm{\theta},\bm{\phi},\bm{\beta})$, and the operator $\cal G$ denotes the channel model (4) that maps the channel parameters $\bm{z}$} to the channel $\bm{H}$}.
@@ III-B1 Operations at UE
where $v_{t}\sim{\cal N}(0,\sigma^{2}_{v})$. Then, the UE sends $r_{t}$ for $t=1,\ldots,T$. Compared to Scheme 1, Scheme 2 bypasses estimating $\bm{H}$ at the UE side. Instead, it directly compresses the received signal and sends it back to the BS. This approach reduces computational overhead and saves time/energy at the BS.
@@ III-B2 MLE Recovery at BS
The compression strategy of Scheme 2 leads to a channel estimation formulation at the BS side that is different from (10). Note that $\tilde{{\bm{y}}}_{t}$ in (11) can be written as
@@
[…]
@@
where $\check{{\bm{v}}}=\check{{\bm{n}}}+{\bm{v}}$ and it can be shown that $\check{{\bm{v}}}\sim{\cal N}(\bm{0},{\mbox{Diag}(\check{\bm{\sigma}}_{v}^{2})})$ for $\check{\bm{\sigma}}_{v}^{2}]_{i}=\sigma_{n}^{2}\sum_{j=1}^{M}a_{i,j}^{2}+{\sigma}_{v}^{2}$ given $\bm{A}$} known as a prior at the BS side. Similar to Scheme 1, the BS estimates $\check{{\bm{h}}}$ through the following ML estimation problem
[…]
@@
$\begin{split}p({\bm{r}}|\check{{\bm{h}}};{{\bm{D}}})=&\leavevmode\nobreak\ \prod_{i=1}^{{T}}p(r_{i}|\check{{\bm{h}}};{{\bm{d}}_{i}})\\ p(r_{i}|\check{{\bm{h}}};{{\bm{d}}_{i}})=&\leavevmode\nobreak\ \Phi\left(\frac{\bar{b}[r_{i}]-{\bm{d}}_{i}^{\top}\check{{\bm{h}}}}{\check{\sigma_{v}}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{d}}_{i}^{\top}\check{{\bm{h}}}}{\check{\sigma_{v}}}\right),\end{split}$
@@
In general, Scheme 1 has more operations performed by the UE as it involves estimating $\bm{H}$}, compression, and quantization. These give rise to $O(MNT+MNR)$ flops. Scheme 2 saves time at UE as it directly compresses the ${\bm{H}}\bm{S}$ without extracting $\bm{H}$} first, which only uses $O(MT)$ flops. Note that the flop reduction at the UE by Scheme 2 can be substantial relative to Scheme 1, as $N$ could be (much) larger than $M$. Such reduction is considered useful for time and energy savings. Nonetheless, as one will see in the next section, Scheme 2 gains in efficiency but loses in sample complexity—that is, Scheme 2 requires slightly more feedback bits to attain the same CSI recovery accuracy of Scheme 1.
@@ IV-A Preliminaries
In the channel model (2), it is reasonable to assume that the channel path loss $\beta_{k}$ is bounded. We denote the maximum amplitude of $\beta_{k}$ as $\kappa$, i.e., $|\beta_{k}|\leq\kappa$ for all ${k}=1,\ldots,K$. The parameter space of $\bm{z}$}, denoted by $\mathcal{Z}$, is then described by
@@
that is, the function $\cal G$} is Lipschitz continuous.
Note that $L_{\cal G}<\infty$ exists because the array manifolds and the double directional channel model do not give rise to erratic changes under continuous inputs $\bm{z}$; see Appendix E in the supplementary material of this paper.
@@
Note that when $x$ is bounded (technically, as $x$ is Gaussian distributed, it does not have a deterministic bounded support, yet the realizations can be assumed to be bounded with a probability of at least $1-\vartheta$ for some $\vartheta\geq 0$), the negative log-likelihood function satisfies
@@
for a certain constant $U_{f}>0$ with probability at least $1-\vartheta$. In addition, as the gradient of $f_{r}(x)$ is continuous, by the extreme value theorem we have, with probability $1-\vartheta$, that
@@
for some constant $F_{f}>0$ with probability $1-\vartheta$. The constant $F_{f}$ being bounded away from zero means that the function $f_{r}(x)$ does not change too slowly.
@@ IV-B Recoverability Analysis of Scheme 1
Assume that the entries of ${\bm{A}}\in\mathbb{R}^{R\times J}$ satisfy $a_{i,j}\sim{\cal N}(0,1/R)$ for all $i,j$, and that $x_{i}$ for all $i\in[R]$ are bounded with probability at least $1-\vartheta$ for $\vartheta\geq 0$. In addition, assume that
@@
[…]
with probability at least $1-3\eta-{\vartheta-}e^{-\Omega(R)}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{R}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}$. and $\psi=\nicefrac{{1}}{{K}}+\nicefrac{{\nu}}{{2}}+2\sqrt{2}+\sqrt{\nicefrac{{2J}}{{R}}}$.
@@ IV-C Recoverability Analysis of Scheme 2
Assume that $\bm{D}$} has a size of $T\times J$ with $J=4MN$, where ${\bm{D}}=\nicefrac{{1}}{{T}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$ that $x_{i}$ for all $i\in[T]$ are bounded with probability at least $1-\vartheta$ for $\vartheta\geq 0$, and that the entries of $\bm{S}$ and $\bm{A}$ are independently drawn from the standard normal distribution. Assume that the following holds:
@@
$\begin{split}&\leavevmode\nobreak\ \frac{\|{\bm{h}}^{\natural}-{\bm{h}}^{\star}\|_{2}^{2}}{MN}\\ \leq&\leavevmode\nobreak\ \frac{64}{\sqrt{T}}\left(4\frac{L_{f}}{F_{f}}(\tau+\frac{3}{4}\nu)+9\frac{U_{f}}{F_{f}}\sqrt{{2}\log(2/\eta)}\right)+\frac{32\psi^{2}}{T}\end{split}$
with probability at least $1-3\eta-{\vartheta-}e^{-\mathbf{\Omega}(\sqrt{T})}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{T}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}$ and $\psi=\nicefrac{{1}}{{K}}+\nicefrac{{\nu}}{{2}}+2\sqrt{2}+\sqrt{\nicefrac{{2J}}{{T}}}$.
The proof of Theorem 2 is relegated to Appendix F in the supplementary material of this paper.